How many cups of party mix that is 64% pretzels must be added to 84 cups of a party mix that is 43% pretzels to make a party that is 50% pretzles?
Answer by stanbon(48545)

How many cups of party mix that is 64% pretzels must be added to 84 cups of a party mix that is 43% pretzels to make a party that is 50% pretzles?
--------------
64% pretzel DATA:
Amt = x cups ; Amt of pretzels = 0.64x cups
-----------
43% pretzel DATA:
Amt = 84 cups ; Amt of pretzels = 0.43*84 = 36.12 cups
-------------
50% pretzel DATA:
Amt = x+84 cups ; Amt of pretzels = 0.50*(x+84)= 0.5x+42 cups
----------
EQUATION:
pret + pret = pret
0.64x + 36.12 = 0.50x+42
0.14x= 5.88
x=42 cups (Amount of 64% pretzel mix that must be added.)
-------------
Cheers,
Stan H.